OpenStudy (anonymous):

Which of the following equals 5^2? 5 10 25 125

OpenStudy (uri):

5*5=25

OpenStudy (anonymous):

When you square a number, you multiply it by itself. So 5^2=5*5=25

OpenStudy (adilalvi):

25

OpenStudy (anonymous):

25